Probability
Get insights from 70 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability
- 70
Questions
- 0
Discussions
- 7
Active Users
- 0
Followers
Have a question about Probability? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
Each element of ordered pair (i, j) is either present in A or in B.
So, A + B = Sum of all elements of all ordered pairs {i, j} for and
= 20 (1 + 2 + 3 + … + 10) = 1100
85. An experiment succeeds twice as often as it fails. Find the probability that in the next six trails, there will be at least 4 successes.
85. An experiment succeeds twice as often as it fails. Find the probability that in the next six trails, there will be at least 4 successes.
85. The probability of success is twice the probability of failure.
Let the probability of failure be x.
∴ Probability of success = 2x
Let and
Let X be the random variable that represents the number of successes in six trials.
By binomial distribution, we obtain
Probability of at lea
75. A person buys a lottery ticket in 50 lotteries, in each of which his chance of winning a prize is What is the probability that he will win a prize:
(a) At least once
(b) Exactly once
(c) At least twice?
75. A person buys a lottery ticket in 50 lotteries, in each of which his chance of winning a prize is What is the probability that he will win a prize:
(a) At least once
(b) Exactly once
(c) At least twice?
75. Let X represent the number of winning prizes in 50 lotteries. The trials are Bernoulli trials.
Clearly, X has a binomial distribution with n = 50 and p = 1/100
P (winning at least once)
P (winning exactly once)
P (at least twice)
16. If P (A) = 1/2 P (B) = 0, then P(A|B) is:
(A) 0
(B) 1/2
(C) Not defined
(D) 1
16. If P (A) = 1/2 P (B) = 0, then P(A|B) is:
(A) 0
(B) 1/2
(C) Not defined
(D) 1
16.
= which is not defined
The correct answer is not defined.
54. The number lock of a suitcase has 4 wheels, each labelled with ten digits i.e., from 0 to 9. The lock opens with a sequence of four digits with no repeats. What is the probability of a Person getting the right sequence to open the suitcase?
54. The number lock of a suitcase has 4 wheels, each labelled with ten digits i.e., from 0 to 9. The lock opens with a sequence of four digits with no repeats. What is the probability of a Person getting the right sequence to open the suitcase?
54. Since, the lock can be open by a combination of four digits from the given ten digits I e, from 0 to 9. The number of ways of selecting 4 digits, = 10C4
This combination of 4 digits can again be arranged within themselves in 41 ways. So, total number of

53. If 4-digit numbers greater than 5,000 are randomly formed from the digits0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, (i) The digits are repeated? (ii) The repetition of digits is not allowed?
53. If 4-digit numbers greater than 5,000 are randomly formed from the digits0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, (i) The digits are repeated? (ii) The repetition of digits is not allowed?
53. (a) No. of ways of forming a four-digit number greater than 5000 from the given digit 0, 1, 3, 5, 7. and digit repetition is allowed can be done in such a way either 5 or 7 and occupy the thousands' place and any of the digits 0, 1, 3, 5, 7 can occupy the remaining 3 places.
Hence,
52. From the employees of a company, 5 persons are selected to represent them inthe managing committee of the company. Particulars of five persons are as follows:
1. No. Name Sex Age in years
2. Harish M 30
3. Rohan M 33
4. Sheetal F 46
5. Alis F 28
6. Salim M 41
A person is selected at random from this group to act as a spokesperson. What is the probability that the spokesperson will be either male or over 35 years?
52. From the employees of a company, 5 persons are selected to represent them inthe managing committee of the company. Particulars of five persons are as follows:
1. No. Name Sex Age in years
2. Harish M 30
3. Rohan M 33
4. Sheetal F 46
5. Alis F 28
6. Salim M 41
A person is selected at random from this group to act as a spokesperson. What is the probability that the spokesperson will be either male or over 35 years?
52. Total no. of person selected to represent the company n (s) = 5.
Let A: person is male.
A = {Harish, Rohan, Salim}
n (A) = 3
And B: person has one 35 yes of age.
B = {Sheetal, Salim}
n (B) = 2
And A ∩ B = {Salim}
n (A ∩ B) = 1
Probability that person is either male or over 35 years.
= P (A ∪ B) = P (A) +
51. A and B are two events such that P(A) = 0.54, P(B) = 0.69 and P(A ∩ B) = 0.35.
Find (i) P(A ∪ B) (ii) P(A´ ∩ B´) (iii) P(A ∩ B´) (iv) P(B ∩ A´)
51. A and B are two events such that P(A) = 0.54, P(B) = 0.69 and P(A ∩ B) = 0.35.
Find (i) P(A ∪ B) (ii) P(A´ ∩ B´) (iii) P(A ∩ B´) (iv) P(B ∩ A´)
51. Given, P (A) = 0.54
P (B) = 0.69.
P (A ∩ B) = 0.35.
(i) P (A ∪ B) = P (A) + P (B) - P (A ∩ B)
= 0.54 + 0.69 - 0.35
= 0. 88
(ii) P (A? ∩ B? ) = P (A ∩ B)? = 1 - P (A ∪ B) = 1 - 0.88 = 0.12
(iii) P (A ∩ B? ) = P (A) - P (A ∩ B)
= 0.54 - 0. 35 = 0.19
(iv) P (B ∩ A? ) = P (B) - P (A ∩ B) = 0.69 - 0.35 =
50. Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.
50. Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.